Find the value of 7 to the -3.
(7 with -3 exponent)

Answer by stanbon(75887)

7^-3
-------
The "negative" means invert.
The 3 means "raise to the 3rd power".
------------
7^-3 = (1/7)^3 = 1/343
==============
Cheers,
Stan H.